OpenStudy (anonymous):

g(x)=6+x+2e^x , find g^-1(8)

OpenStudy (anonymous):

that means to find the inverse of g(8)

OpenStudy (anonymous):

I understand that i need to find the inverse, i just don't know what I need to do with the "e"

OpenStudy (anonymous):

err...im not sure about that..here let @Hero help you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

everything is moved to the bottom. 5^-1 = 1/5 1/[6 + 8 + 2e^8] what math are you taking?

OpenStudy (anonymous):

Calculus I

OpenStudy (anonymous):

@myfacewhen that only goes for numbers....but when it says \(g^{-1}\) its not the same as \({1\over g}\)

hero (hero):

If I help, I'm just going to present solution steps. I'm not exactly interested in "teaching" at this point in time. I think @smoothmath might be the guy if you want step by step "guidance".

OpenStudy (anonymous):

Better just go ahead, Hero. I'm unfortunately blanking on how to explain this.

OpenStudy (anonymous):

If i see the answer, I can probably work the problem backwards to figure out where I'm getting stuck.

OpenStudy (anonymous):

Try alpha if you just need the answer. Type "inverse function of"

OpenStudy (anonymous):

lol brain fart yea sorry bout that y = 6 + x + 2e^x ln e^x = x swap x for y x = 6 + y + 2e^y solve for y I can actually solve it if you need me to sorry it took so long I had some private business to attend to.

OpenStudy (anonymous):

I usually can solve for y but the e is what is throwing me off

OpenStudy (anonymous):

e is a constant

OpenStudy (anonymous):

treat it like you would treat \(a\) or \(x\) or \(\pi\)

OpenStudy (anonymous):

I believe the e has to do with the exponential and logarithmic functions that we are working on in class. I think I cancel out the e with natural log but I'm not getting the correct answer

OpenStudy (zarkon):

this problem can be easily solved by inspection

hero (hero):

You can't replace y with 8

hero (hero):

However, you can do this: y = 6 + 8 + 2e^8 Then solve for y Once you have the point, you can reflect it over the axis y = x to get the inverse

OpenStudy (anonymous):

thank you.

OpenStudy (zarkon):

\(g(x)=6+x+2e^x\) notice \(g(0)=6+0+2e^0=6+2=8\) therefore \(0=g^{-1}(8)\)